# Regularity Propagation of Global Weak Solutions to a Navier–Stokes–Cahn–Hilliard System for Incompressible Two-phase Flows with Chemotaxis and Active Transport

### v1

@@ 4.2 The $\omega$-limit set and eventual separation property of $\varphi$
Below we show that $(\varphi_{\infty},\sigma_{\infty})\in(W^{2,6}(\Omega)\cap H^{2}_{N}(\Omega))\times(W^{2,6}(\Omega)\cap H^{2}_{N}(\Omega))$ is a strong solution to the stationary problem (2.4a)–(2.4c). The boundary condition (2.4c) is obvious. Besides, (2.4b) is a consequence of (4.11), that is, $\sigma_{\infty}-\chi\varphi_{\infty}$ is indeed a constant. Thanks to the estimate (4.4) on $\partial_{t}\varphi$ and $\partial_{t}\sigma$, by an argument similar to that in [28, Section 5] (see also [33, Section 6], [45, Section 4]), we can verify (2.4a).
Concerning the strict separation property of $\varphi_{\infty}$, we apply a dynamic approach inspired by [41]. The key observation is that every $(\varphi_{\infty},\sigma_{\infty})$ can be viewed as (at least) a global weak solution to the evolution problem (3.1)–(3.3) with $\gamma\in(0,1)$, $\bm{v}=\mathbf{0}$ and the initial datum given by $(\varphi_{\infty},\sigma_{\infty})$ itself. Then it follows from [28, Proposition 5.1] that (4.6) holds with some $\delta_{0}\in(0,1)$. Moreover, since the set $\{\varphi_{\infty}\}$ is compact in $W^{2-\epsilon,6}(\Omega)\hookrightarrow C(\overline{\Omega})$ for $\epsilon\in(0,1/2)$, we can find $\delta_{0}\in(0,1)$ independent of $\varphi_{\infty}$ such that (4.6) holds (cf. [4]). The proof of Proposition 4.1 is complete. ∎
@@
Define $\omega(\varphi)=\{\varphi_{\infty}:(\mathbf{0},\varphi_{\infty},\sigma_{\infty})\in\omega(\bm{v},\varphi,\sigma)\}$. It follows from the relative compactness of the trajectory $\varphi(t)$ and Proposition 4.1 that
@@ 4.3 Weak-strong uniqueness and eventual regularity of $(\bm{v},\sigma)$
To achieve the eventual regularity of the velocity field $\bm{v}$, like for the classical Navier–Stokes equations in three dimensions, we need a weak-strong uniqueness result for the full system (1.1)–(1.3).
@@
Let $\Omega$ be a bounded domain in $\mathbb{R}^{3}$, with boundary $\partial\Omega$ of class $C^{3}$. Suppose that the hypotheses (H1)–(H3) are satisfied. Consider the initial datum $(\bm{v}_{0},\varphi_{0},\sigma_{0})$ satisfing $\bm{v}_{0}\in\bm{H}^{1}_{0,\mathrm{div}}(\Omega)$, $\varphi_{0}\in H^{2}_{N}(\Omega)$, $\|\varphi_{0}\|_{L^{\infty}}\leq 1$, $|\overline{\varphi_{0}}|<1$, $\mu_{0}=-\Delta\varphi_{0}+\varPsi^{\prime}(\varphi_{0})\in H^{1}(\Omega)$ and $\sigma_{0}\in H^{1}(\Omega)$. Furthermore, we suppose that:
@@
$\bm{v}_{1}\in L^{\infty}(0,T;\bm{L}^{2}_{0,\mathrm{div}}(\Omega))\cap L^{2}(0,T;\bm{H}^{1}_{0,\mathrm{div}}(\Omega))\cap W^{1,\frac{4}{3}}(0,T;\bm{H}^{1}_{0,\mathrm{div}}(\Omega)^{\prime}),$
$\varphi_{1}\in L^{\infty}(0,T;W^{2,6}(\Omega)\cap H^{2}(\Omega)),\quad\partial_{t}\varphi_{1}\in L^{\infty}(0,T;(H^{1}(\Omega))^{\prime})\cap L^{2}(0,T;H^{1}(\Omega)),$
@@
$\sigma_{1}\in L^{\infty}(0,T;L^{6}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega)),\quad\partial_{t}\sigma_{1}\in L^{2}(0,T;H^{1}(\Omega)^{\prime}),$
@@
$\left\langle\partial_{t}\bm{v}_{1},\bm{\zeta}\right\rangle_{(\bm{H}^{1}_{0,\mathrm{div}})^{\prime},\bm{H}^{1}_{0,\mathrm{div}}}+((\bm{v}_{1}\cdot\nabla)\bm{v}_{1},\bm{\zeta})+(2\nu(\varphi_{1})D\bm{v}_{1},D\bm{\zeta})=((\mu_{1}+\chi\sigma_{1})\nabla\varphi_{1},\bm{\zeta}),$ (4.13a)
$\left\langle\partial_{t}\sigma_{1},\xi\right\rangle_{(H^{1})^{\prime},H^{1}}+({\bm{v}_{1}\cdot\nabla\sigma_{1}},\xi)+(\nabla\sigma_{1},\nabla\xi)=\chi(\nabla\varphi_{1},\nabla\xi),$ (4.13b)
for any $\bm{\zeta}\in\bm{H}^{1}_{0,\mathrm{div}}$, $\xi\in H^{1}(\Omega)$ and almost all $t\in(0,T)$,
$\partial_{t}\varphi_{1}+\bm{v}_{1}\cdot\nabla\varphi_{1}=\Delta\mu_{1}-\alpha(\overline{\varphi_{1}}-c_{0}),$ (4.14a)
$\mu_{1}=-\Delta\varphi_{1}+\varPsi^{\prime}(\varphi_{1})-\chi\sigma_{1}+\beta\mathcal{N}(\varphi_{1}-\overline{\varphi_{1}}),$ (4.14b)
@@
$\sigma_{2}\in L^{\infty}(0,T;L^{6}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega)),\quad\partial_{t}\sigma_{2}\in L^{2}(0,T;H^{1}(\Omega)^{\prime}).$
The solution $(\bm{v}_{2},\varphi_{2},\mu_{2},\sigma_{2})$ fulfills the equations (1.1)–(1.1d) almost everywhere in $\Omega\times(0,T)$, and
$\left\langle\partial_{t}\sigma_{2},\xi\right\rangle_{(H^{1})^{\prime},H^{1}}+({\bm{v}_{2}\cdot\nabla\sigma_{2}},\xi)+(\nabla\sigma_{2},\nabla\xi)=\chi(\nabla\varphi_{2},\nabla\xi),$ (4.15a)
@@
Then we have $(\bm{v}_{1},\varphi_{1},\mu_{1},\sigma_{1})=(\bm{v}_{2},\varphi_{2},\mu_{2},\sigma_{2})$ on $[0,T]$.
@@
Denote the differences of solutions by
@@
Then by assumption, it holds $\bm{v}(0)=\mathbf{0}$, $\varphi(0)=\sigma(0)=0$, and $\overline{\varphi(t)}=\overline{\sigma(t)}=0$ for all $t\in[0,T]$. Moreover, there exists $\delta_{*}\in(0,1)$ such that
@@
The strict separation property of $\varphi_{1}$, $\varphi_{2}$ will play an crucial role in the subsequent proof.
The weak solution $(\bm{v}_{1},\varphi_{1},\mu_{1},\sigma_{1})$ satisfies the energy inequality (2.3) on $[0,T]$. Since $(\bm{v}_{2},\varphi_{2},\mu_{2},\sigma_{2})$ is sufficiently regular, it satisfies the corresponding energy equality. Besides, we see that $\bm{v}_{1}$ can be used as a test function for the equation of $\bm{v}_{2}$ and vice versa. Adding the resultants together and integrating on $[0,t]\subset[0,T]$, we obtain
@@
Subtracting (4.16) from the sum of the energy inequality (2.3) for $(\bm{v}_{1},\varphi_{1},\mu_{1},\sigma_{1})$ with the corresponding energy equality for $(\bm{v}_{2},\varphi_{2},\mu_{2},\sigma_{2})$, using the incompressibility condition and integration by parts, we find
@@
On the other hand, thanks to Theorem 3.2, both solutions satisfies the energy equality (3.8) on $[0,T]$. Subtracting them from the inequality (4.17) yields
@@
Testing the equation for the difference $\varphi$ by $-\Delta\varphi$, integrating over $\Omega\times[0,t]$ gives
@@
Since $\mathcal{Z}\in L^{1}(0,T)$, we infer from (4.22) and Gronwall’s lemma that $\mathcal{Y}(t)=0$ for all $t\in[0,T]$. This combined with the Poincaré–Wirtinger inequality leads to the desired conclusion.
@@
[…]
for given data $c$, $\bm{u}_{0}$, $\bm{f}$. Suppose that the hypotheses (H1) is satisfied, $c\in BUC([0,+\infty);W^{1,6}(\Omega))$, $\bm{u}_{0}\in\bm{H}^{1}_{0,\mathrm{div}}(\Omega)$ and $\bm{f}\in L^{2}(0,+\infty;\bm{L}^{2}_{0,\mathrm{div}}(\Omega))$. There is some $\varepsilon_{0}>0$ such that, if
$\|\bm{u}_{0}\|_{\bm{H}^{1}_{0,\mathrm{div}}(\Omega)}+\|\bm{f}\|_{L^{2}(0,+\infty;\bm{L}^{2}_{0,\mathrm{div}}(\Omega))}\leq\varepsilon_{0},$
then there is a unique solution $\bm{u}\in L^{2}(0,T;D(\bm{S}))\cap L^{2}(0,T;\bm{L}^{2}_{0,\mathrm{div}}(\Omega))$ of problem (4.23)–(4.26) with $T=+\infty$.
@@
Now taking $(\bm{v}(T_{\mathrm{R}}),\varphi(T_{\mathrm{R}}),\sigma(T_{\mathrm{R}}))$ as the initial datum, according to [30, Theorem 3.1], problem (1.1)–(1.3) admits a unique local strong solution $(\widehat{\bm{v}},\widehat{p},\widehat{\varphi},\widehat{\mu},\widehat{\sigma})$ defined on the maximal interval $\left[T_{\mathrm{R}},T_{\max}\right)$ such that for any $\widetilde{T}\in[T_{\mathrm{R}},T_{\max})$, it fulfills
@@
$\widehat{\varphi}\in L^{\infty}(T_{\mathrm{R}},\widetilde{T};W^{2,6}(\Omega)),\quad\partial_{t}\varphi\in L^{\infty}(T_{\mathrm{R}},\widetilde{T};(H^{1}(\Omega))^{\prime})\cap L^{2}(T_{\mathrm{R}},\widetilde{T};H^{1}(\Omega)),$
@@
$\widehat{\mu}\in L^{\infty}(T_{\mathrm{R}},\widetilde{T};H^{1}(\Omega))\cap L^{2}(T_{\mathrm{R}},\widetilde{T};H^{3}(\Omega)),$
$\widehat{\sigma}\in C([T_{\mathrm{R}},\widetilde{T}];H^{1}(\Omega))\cap L^{2}(T_{\mathrm{R}},\widetilde{T};H^{2}(\Omega))\cap H^{1}(T_{\mathrm{R}},\widetilde{T};L^{2}(\Omega)),$
@@
The solution $(\widehat{\bm{v}},\widehat{p},\widehat{\varphi},\widehat{\mu},\widehat{\sigma})$ satisfies the system (1.1) almost everywhere in $\Omega\times(T_{\mathrm{R}},T_{\max})$, the boundary condition (1.2) almost everywhere on $\partial\Omega\times(T_{\mathrm{R}},T_{\max})$, and the initial conditions $\widehat{\bm{v}}(T_{\mathrm{R}})=\bm{v}(T_{\mathrm{R}})$, $\widehat{\varphi}(T_{\mathrm{R}})=\varphi(T_{\mathrm{R}})$, $\widehat{\sigma}(T_{\mathrm{R}})=\sigma(T_{\mathrm{R}})$ almost everywhere in $\Omega$. Moreover, it satisfies the energy equality:
@@
for all $T_{\mathrm{R}}\leq t_{1}\leq t_{2}\leq\widetilde{T}$. Then we can apply Theorem 2.1-(1), (2) and Proposition 4.2 to conclude that
@@
As a consequence, in order to complete the proof of Theorem 2.1-(3), it remains to show that $T_{\max}=+\infty$. To this end, we notice that (4.3), (4.4) and the separation property (4.12) hold globally in time, and they are independent of $T_{\max}$. Thanks to (4.27), below we will not distinguish $(\bm{v},\varphi,\mu,\sigma)$, $(\widehat{\bm{v}},\widehat{\varphi},\widehat{\mu},\widehat{\sigma})$ since they coincide as long as the later exists.
@@
with $\bm{f}=-\varphi\nabla\mu-\chi\varphi\nabla(\sigma-\chi\varphi)$ and $\nabla z=\bm{f}-\bm{g}$. Then it follows from (4.1), (4.3) and the property of the Leray projection $\bm{P}$ that
@@
Since $\varphi\in BUC([T_{\mathrm{R}},+\infty);W^{1,6}(\Omega))$, $\|\bm{v}(T_{\mathrm{R}})\|_{\bm{H}^{1}_{0,\mathrm{div}}(\Omega))}+\|\bm{g}\|_{L^{2}(T_{\mathrm{R}},+\infty;\bm{L}^{2}_{0,\mathrm{div}}(\Omega))}\leq\varepsilon_{0}$, then we can apply Lemma 4.1 to conclude that the equation (4.28) subject to the no-slip boundary condition and the initial condition $\bm{v}(T_{\mathrm{R}})$ admits a unique global strong solution $\bm{v}\in L^{2}(T_{\mathrm{R}},+\infty;D(\bm{S}))\cap H^{1}(T_{\mathrm{R}},+\infty;\bm{L}^{2}_{0,\mathrm{div}}(\Omega))$. By interpolation, we also have $\bm{v}\in BUC([T_{\mathrm{R}},+\infty);\bm{H}^{1}_{0,\mathrm{div}}(\Omega))$.
Next, testing the equation (1.1e) for $\sigma$ by $-\Delta\sigma$ and integrating over $\Omega$, we get
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\nabla\sigma\|^{2}+\|\Delta\sigma\|^{2}=(\bm{v}\cdot\nabla\sigma,\Delta\sigma)+\chi(\Delta\varphi,\Delta\sigma).$
By a similar argument for (3.62) and the continuous embedding $\bm{H}^{2}(\Omega)\hookrightarrow\bm{L}^{\infty}(\Omega)$, we find that
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\nabla\sigma\|^{2}+\frac{1}{2}\|\Delta\sigma\|^{2}\leq C\|\bm{v}\|_{\bm{H}^{2}}^{2}\|\nabla\sigma\|^{2}+\chi^{2}\|\Delta\varphi\|^{2}.$
Since $\sigma(T_{\mathrm{R}})\in H^{1}(\Omega)$, $\bm{v}\in L^{2}(T_{\mathrm{R}},+\infty;D(\bm{S}))$ and $\varphi\in L^{4}_{\mathrm{uloc}}(T_{\mathrm{R}},+\infty;H^{2}(\Omega))$ (thanks to (4.3) and a similar estimate for (3.30)), we can apply the classical Gronwall lemma on $[T_{\mathrm{R}},T_{\mathrm{R}}+1]$ and the uniform Gronwall lemma on $[T_{\mathrm{R}}+1,+\infty)$ to conclude that

### v2

@@ 4.2 The $\omega$-limit set and eventual separation property of $\varphi$
Below we show that $(\varphi_{\infty},\sigma_{\infty})\in(W^{2,6}(\Omega)\cap H^{2}_{N}(\Omega))\times(W^{2,6}(\Omega)\cap H^{2}_{N}(\Omega))$ is a strong solution to the stationary problem (2.6a)–(2.6c). The boundary condition (2.6c) is obvious. Besides, (2.6b) is a consequence of (4.12), that is, $\sigma_{\infty}-\chi\varphi_{\infty}$ is indeed a constant. Thanks to the estimate (4.4) on $\partial_{t}\varphi$ and $\partial_{t}\sigma$, by an argument similar to that in [32, Section 5] (see also [38, Section 6], [53, Section 4]), we can verify (2.6a).
Concerning the strict separation property of $\varphi_{\infty}$, we apply a dynamic approach inspired by [47]. The key observation is that every $(\varphi_{\infty},\sigma_{\infty})$ can be viewed as (at least) a global weak solution to the evolution problem (3.1)–(3.3) with $\gamma\in(0,1)$, $\bm{v}=\mathbf{0}$ and the initial datum given by $(\varphi_{\infty},\sigma_{\infty})$ itself. Then it follows from [32, Proposition 5.1] that (4.7) holds with some $\delta_{0}\in(0,1)$. Moreover, since the set $\{\varphi_{\infty}\}$ is compact in $W^{2-\epsilon,6}(\Omega)\hookrightarrow C(\overline{\Omega})$ for $\epsilon\in(0,1/2)$, we can find $\delta_{0}\in(0,1)$ independent of $\varphi_{\infty}$ such that (4.7) holds (cf. [5]). The proof of Proposition 4.1 is complete. ∎
@@
Define $\omega(\varphi)=\{\varphi_{\infty}:(\mathbf{0},\varphi_{\infty},\sigma_{\infty})\in\omega(\bm{v},\varphi,\sigma)\}$. It follows from the relative compactness of the trajectory $\varphi(t)$ in $W^{2-\epsilon,6}(\Omega)$ and Proposition 4.1 that
@@ 4.3 Weak-strong uniqueness and eventual regularity of $(\bm{v},\sigma)$
To establish the eventual regularity of the velocity field $\bm{v}$, as for the classical Navier–Stokes equations in three dimensions, we need a weak-strong uniqueness result for the full system (1.1)–(1.3).
@@
Let $\Omega$ be a bounded domain in $\mathbb{R}^{3}$, with boundary $\partial\Omega$ of class $C^{3}$. Suppose that hypotheses (H1)–(H3) are satisfied. Consider the initial datum $(\bm{v}_{0},\varphi_{0},\sigma_{0})$ that satisfies $\bm{v}_{0}\in\bm{H}^{1}_{0,\mathrm{div}}(\Omega)$, $\varphi_{0}\in H^{2}_{N}(\Omega)$, $\|\varphi_{0}\|_{L^{\infty}}\leq 1$, $|\overline{\varphi_{0}}|<1$, $\mu_{0}=-\Delta\varphi_{0}+\varPsi^{\prime}(\varphi_{0})\in H^{1}(\Omega)$ and $\sigma_{0}\in H^{1}(\Omega)$. Furthermore, we suppose that:
@@
$\bm{v}_{1}\in L^{\infty}(0,T;\bm{L}^{2}_{0,\mathrm{div}}(\Omega))\cap L^{2}(0,T;\bm{H}^{1}_{0,\mathrm{div}}(\Omega))\cap W^{1,\frac{4}{3}}(0,T;(\bm{H}^{1}_{0,\mathrm{div}}(\Omega))^{\prime}),$
$\varphi_{1}\in L^{\infty}(0,T;W^{2,6}(\Omega)\cap H^{2}_{N}(\Omega)),\quad\partial_{t}\varphi_{1}\in L^{\infty}(0,T;(H^{1}(\Omega))^{\prime})\cap L^{2}(0,T;H^{1}(\Omega)),$
@@
$\sigma_{1}\in L^{\infty}(0,T;L^{6}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega)),\quad\partial_{t}\sigma_{1}\in L^{2}(0,T;(H^{1}(\Omega))^{\prime}),$
@@
$\left\langle\partial_{t}\bm{v}_{1},\bm{\zeta}\right\rangle_{(\bm{H}^{1}_{0,\mathrm{div}})^{\prime},\bm{H}^{1}_{0,\mathrm{div}}}+((\bm{v}_{1}\cdot\nabla)\bm{v}_{1},\bm{\zeta})+(2\nu(\varphi_{1})D\bm{v}_{1},D\bm{\zeta})=((\mu_{1}+\chi\sigma_{1})\nabla\varphi_{1},\bm{\zeta}),$
$\left\langle\partial_{t}\sigma_{1},\xi\right\rangle_{(H^{1})^{\prime},H^{1}}+({\bm{v}_{1}\cdot\nabla\sigma_{1}},\xi)+(\nabla\sigma_{1},\nabla\xi)=\chi(\nabla\varphi_{1},\nabla\xi),$
for any $\bm{\zeta}\in\bm{H}^{1}_{0,\mathrm{div}}(\Omega)$, $\xi\in H^{1}(\Omega)$ and almost all $t\in(0,T)$,
$\partial_{t}\varphi_{1}+\bm{v}_{1}\cdot\nabla\varphi_{1}=\Delta\mu_{1}-\alpha(\overline{\varphi_{1}}-c_{0}),$
$\mu_{1}=-\Delta\varphi_{1}+\varPsi^{\prime}(\varphi_{1})-\chi\sigma_{1}+\beta\mathcal{N}(\varphi_{1}-\overline{\varphi_{1}}),$
@@
$\sigma_{2}\in L^{\infty}(0,T;L^{6}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega)),\quad\partial_{t}\sigma_{2}\in L^{2}(0,T;(H^{1}(\Omega))^{\prime}).$
The solution $(\bm{v}_{2},\varphi_{2},\mu_{2},\sigma_{2})$ fulfills the equations (1.1a)–(1.1d) almost everywhere in $\Omega\times(0,T)$, and
$\left\langle\partial_{t}\sigma_{2},\xi\right\rangle_{(H^{1})^{\prime},H^{1}}+({\bm{v}_{2}\cdot\nabla\sigma_{2}},\xi)+(\nabla\sigma_{2},\nabla\xi)=\chi(\nabla\varphi_{2},\nabla\xi),$
@@
Then we have
$(\bm{v}_{1},\varphi_{1},\mu_{1},\sigma_{1})=(\bm{v}_{2},\varphi_{2},\mu_{2},\sigma_{2})\quad\text{on}\ \ [0,T].$
@@
Denote the differences of two solutions by
@@
By assumption, we have
$\bm{v}(0)=\mathbf{0},\quad\varphi(0)=\sigma(0)=0,\quad\overline{\varphi(t)}=\overline{\sigma(t)}=0,\quad\forall\,t\in[0,T].$
Moreover, there exists $\delta_{*}\in(0,1)$ such that
@@
The strict separation property of $\varphi_{1}$, $\varphi_{2}$ will play a crucial role in the subsequent proof.
The solution $(\bm{v}_{1},\varphi_{1},\mu_{1},\sigma_{1})$ satisfies the energy inequality (2.5) on $[0,T]$. Since $(\bm{v}_{2},\varphi_{2},\mu_{2},\sigma_{2})$ is sufficiently regular, it satisfies the corresponding energy equality. In addition, we see that $\bm{v}_{1}$ can be used as a test function for the equation of $\bm{v}_{2}$ and vice versa. Adding the resultants together and integrating on $0,t]\subset[0,T]$, we obtain
@@
Subtracting (4.16) from the sum of the energy inequality (2.5) for $(\bm{v}_{1},\varphi_{1},\mu_{1},\sigma_{1})$ with the corresponding energy equality for $(\bm{v}_{2},\varphi_{2},\mu_{2},\sigma_{2})$, using the incompressibility condition for $\bm{v}_{1}$, $\bm{v}_{2}$, and integration by parts, we find
@@
On the other hand, due to Theorem 3.2, both solutions $(\varphi_{1},\mu_{1},\sigma_{1})$, $(\varphi_{2},\mu_{2},\sigma_{2})$ satisfy the energy equality (3.7) on $[0,T]$ with $\bm{v}=\bm{v}_{1},\bm{v}_{2}$, respectively. Subtracting them from the inequality (4.17) yields
@@
Testing the equation for the difference $\varphi$ by $\Delta\varphi$, integrating over $\Omega\times[0,t]$ gives
@@
Since $\mathcal{Z}\in L^{1}(0,T)$, we infer from (4.22) and Gronwall’s lemma that
$\mathcal{Y}(t)=0,\quad\forall\,t\in[0,T].$
This combined with the Poincaré–Wirtinger inequality leads to the desired conclusion.
@@
$\partial_{t}\bm{u}+\bm{u}\cdot\nabla\bm{u}-\mathrm{div}\big(2\nu(c)D\bm{u}\big)+\nabla p=\bm{f},$ $\text{ in }\Omega\times(0,T),$ (4.23)
$\mathrm{div}\,\bm{u}=0,$ $\text{ in }\Omega\times(0,T),$ (4.24)
$\bm{u}=\mathbf{0},$ $\text{ on }\partial\Omega\times(0,T),$ (4.25)
$\bm{u}|_{t=0}=\bm{u}_{0},$ $\text{ in }\Omega,$ (4.26)
for given data $c$, $\bm{u}_{0}$, $\bm{f}$. Suppose that the assumption (H1) is satisfied, $c\in BUC([0,+\infty);W^{1,6}(\Omega))$, $\bm{u}_{0}\in\bm{H}^{1}_{0,\mathrm{div}}(\Omega)$ and $\bm{f}\in L^{2}(0,+\infty;\bm{L}^{2}_{0,\mathrm{div}}(\Omega))$. There is some $\varepsilon_{0}>0$ such that, if
$\|\bm{u}_{0}\|_{\bm{H}^{1}_{0,\mathrm{div}}}+\|\bm{f}\|_{L^{2}(0,+\infty;\bm{L}^{2}_{0,\mathrm{div}}(\Omega))}\leq\varepsilon_{0},$
then there is a unique solution $\bm{u}\in L^{2}(0,T;D(\bm{S}))\cap H^{1}(0,T;\bm{L}^{2}_{0,\mathrm{div}}(\Omega))$ of problem (4.23)–(4.26) with $T=+\infty$.
@@
[…]
@@
$\widehat{\varphi}\in L^{\infty}(T_{\mathrm{R}},\widetilde{T};W^{2,6}(\Omega)\cap H^{2}_{N}(\Omega)),\quad\partial_{t}\varphi\in L^{\infty}(T_{\mathrm{R}},\widetilde{T};(H^{1}(\Omega))^{\prime})\cap L^{2}(T_{\mathrm{R}},\widetilde{T};H^{1}(\Omega)),$
@@
$\widehat{\mu}\in L^{\infty}(T_{\mathrm{R}},\widetilde{T};H^{1}(\Omega))\cap L^{2}(T_{\mathrm{R}},\widetilde{T};H^{3}(\Omega)\cap H^{2}_{N}(\Omega)),$
$\widehat{\sigma}\in C([T_{\mathrm{R}},\widetilde{T}];H^{1}(\Omega))\cap L^{2}(T_{\mathrm{R}},\widetilde{T};H^{2}_{N}(\Omega))\cap H^{1}(T_{\mathrm{R}},\widetilde{T};L^{2}(\Omega)),$
@@
[…]
@@
for all $T_{\mathrm{R}}\leq t_{1}\leq t_{2}\leq\widetilde{T}$. Thus, we can apply Theorem 2.1-(1), (2) and Proposition 4.2 to conclude that
@@
[…]
@@
and
$\bm{f}=-\varphi\nabla\mu-\chi\varphi\nabla(\sigma-\chi\varphi),\quad\nabla z=\bm{f}-\bm{g}.$
Then it follows from (4.1), (4.3) and the property of the Leray projection $\bm{P}$ that
@@
Since
$\varphi\in BUC([T_{\mathrm{R}},+\infty);W^{1,6}(\Omega)),\quad\|\bm{v}(T_{\mathrm{R}})\|_{\bm{H}^{1}_{0,\mathrm{div}}(\Omega))}+\|\bm{g}\|_{L^{2}(T_{\mathrm{R}},+\infty;\bm{L}^{2}_{0,\mathrm{div}}(\Omega))}\leq\varepsilon_{0},$
then we can apply Lemma 4.1 to conclude that equation (4.28) subject to the no-slip boundary condition and the initial condition $\bm{v}(T_{\mathrm{R}})$ admits a unique global strong solution such that
$\bm{v}\in L^{2}(T_{\mathrm{R}},+\infty;D(\bm{S}))\cap H^{1}(T_{\mathrm{R}},+\infty;\bm{L}^{2}_{0,\mathrm{div}}(\Omega)).$
By interpolation, we also find $\bm{v}\in BUC([T_{\mathrm{R}},+\infty);\bm{H}^{1}_{0,\mathrm{div}}(\Omega))$.
Next, testing equation (1.1e) for $\sigma$ by $\Delta\sigma$ and integrating over $\Omega$, we get
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\nabla\sigma\|^{2}+\|\Delta\sigma\|^{2}$ $=(\bm{v}\cdot\nabla\sigma,\Delta\sigma)+\chi(\Delta\varphi,\Delta\sigma)$
$\leq\|\bm{v}\|_{\bm{L}^{\infty}}\|\nabla\sigma\|\|\Delta\sigma\|+|\chi|\|\Delta\varphi\|\|\Delta\sigma\|$
$\leq\frac{1}{2}\|\Delta\sigma\|^{2}+\|\bm{v}^{k}\|_{\bm{L}^{\infty}}^{2}\|\nabla\sigma\|^{2}+\chi^{2}\|\Delta\varphi\|^{2}.$
Since $\sigma(T_{\mathrm{R}})\in H^{1}(\Omega)$, $\bm{v}\in L^{2}(T_{\mathrm{R}},+\infty;D(\bm{S}))$ and $\varphi\in L^{4}_{\mathrm{uloc}}([T_{\mathrm{R}},+\infty);H^{2}(\Omega))$ (thanks to (4.3) and a similar estimate for (3.28)), keeping in mind the continuous embedding $\bm{H}^{2}(\Omega)\hookrightarrow\bm{L}^{\infty}(\Omega)$, we can apply the classical Gronwall’s lemma on $T_{\mathrm{R}},T_{\mathrm{R}}+1]$ and the uniform Gronwall lemma on $T_{\mathrm{R}}+1,+\infty)$ to conclude that